# From Sparse Dependence to Sparse Attention: Unveiling How Chain-of-Thought Enhances Transformer Sample Efficiency

### v1

@@ 1 Introduction
    Theoretically, we show that while the parity problem can be expressed by a 1-layer transformer, learning requires exponential to $k$ samples without CoT when the number of parameters is limited (Theorem 1 and Theorem 2). Meanwhile, for CoT data with sparse sequential dependency, a 1-layer transformer can learn the parity function and faithfully represent the sparse dependency in its attention pattern with almost linear samples with respect to $k$ and $n$ (Theorem 3).
@@ 3.1 Notations and Definitions
Parity Problem. Each token in our setting is either 00 or $1$. We represent a sequence of length $T$ as $\textbf{b}[1],\ldots,\textbf{b}[T]$. A parity function is a function of form $\mathrm{parity}_{S}(\textbf{b})=\oplus_{j\in\mathcal{S}}\textbf{b}[j]$, where $\mathcal{S}$ is the set of secret indices and $\oplus$ is the XOR operator on binary variables. The cardinal of the secret set $k=|S|$ controls the hardness of the problem.
@@
Loss function. To simplify our analysis, we use hinge loss $\ell(\hat{y},y)=\max\{(-1)^{y}\hat{y}+1,0\}$ as the loss function. Consequently, we define the next token prediction loss of a boolean sequence b as $L(w)=\sum_{i=n+1}^{n+k}\ell(\mathcal{T}^{(L)}(\textbf{b})[i],\textbf{b}[i+1])$, where $w$ denotes all the trainable parameters $[A,W]$ in the attention module and the FFN map across all layers.
@@ 3.2 Exponential Sample Complexity without CoT
Consider the Transformer model defined in Definition 6, for any $\delta<0.1$ and large enough $n$, with probability at least $1-\delta$ over the randomness of embedding $e$, there exists a weight configuration of the Transformer with dimension $d=\tilde{\Omega}(k)$ and width $2m=\Omega(k)$, such that it can achieve perfect accuracy on the parity problem $(n,k)$ without CoT, i.e.
$\forall\textbf{b}\sim p_{\rm{NoCoT}},\mathrm{sgn}\left(\mathcal{T}(\textbf{b})[n+1]\right)=\textbf{b}[n+2].$
In other words, the model possesses sufficient expressive power to represent any parity function, even in the absence of the chain of thought. Therefore, we are in the representational-sufficient regime where expressiveness is not the bottleneck. The proof of this statement is based on random matrix theory and concentration inequalities, which we defer to Section A.2.
@@
The proof is deferred to Section A.3. We use results from the classical online learning communities [40] to show that in the regime where the memory required to perform the training is less than $nk$, the model can’t effectively store information about enough samples inside the parameters during training, and hence can’t infer the secret indices effectively. Here the parameters $k$ denotes the size of the secret set $\mathcal{S}$ and the sample complexity grows exponentially when $k$ grows.
@@ 3.3 Polynomial Sample Complexity with CoT
We will use the following initialization of the 1-layer Transformer model. The attention is initialized to be uniform ($A=0$) and the contribution of attention output is initialized to be zero ($W_{r,1:m}=0$). The parameter $h$ in the FFN and the word embedding $\mathbf{E}$ is fixed during training.
@@
Furthermore, after training, the attention pattern is interpretable and one-hot in the sense that, for any $i\in[n+1,n+k]$,
@@
This theorem indicates that the attention module can successfully extract the sparse sequential dependencies in the CoT data and faithfully represent it in the attention pattern, which is validated in our experiments (see Figure 3). It is also amongst the first optimization dynamics analysis of Transformers using finite-sample gradients rather than population gradients.
@@
Phase 1. Configuring the FFN At initialization, as the FFN weight corresponds to the attention output ($W_{r,d+1:2d}$) is initialized to be zero. The set of neurons activated at each position $i\in[n+1,n+k]$ is solely determined by the word embedding $e_{(i,\textbf{b}[i])}$ at the position. Hence, at initialization, the MLP can be viewed as an ensemble of multiple linear functions specialized to each position and boolean value. We can show that when the learning rate is small, the set of activated neurons at each position remains the same through the training process. Hence, we can conceptually view the FFN function as a set of different linear functions applied independently at each position and value. We will denote this set of linear weight as $\kappa_{t,i,b}$ (formally defined in Lemma 18).
@@ 4 Empirical Experiments
In this section, we validate and extend our theoretical findings through comprehensive experiments in the following three aspects: First, we empirically confirm that CoT reduces the sample complexity of transformers in learning the parity problem. Second, we conduct multi-pass training of transformers without CoT, as a complement to the established lower bound (Theorem 2) which only applies to one-pass training. The results indicate that multi-pass training indeed improves the models’ ability to learn parity problems. Third, we conduct experiments on the GSM8K dataset [12] to show that CoT introduces sparse sequential dependence on real-world training data. Although necessary simplifications on transformer are made in Section 3 to develop theoretical results, we use the standard transformer architecture in the subsequent experiments following the GPT-2 architecture [50] with trainable position embeddings unless otherwise specified.
@@ 4.3 CoT Induces Sparsity on Real-World Data
Experiment protocols. In Figure 6, we examine two data types: With CoT, where inputs from GSM8K dataset are concatenated with ground truth answers that include multiple reasoning steps, and No CoT, where inputs are directly concatenated with the final answer. We evaluate two language models: the pre-trained model Qwen2-7B [68] and the specialized mathematics model Qwen2-7B-Math [49] which is fine-tuned from Qwen2-7B on a mathematics-specific corpus with CoT data. We plot the normalized attention entropy (Equation 1) across different heads. More details can be found in the Appendix Section B.1.
@@ 5 Additional Related Work
Parity Learning. The most relevant work to ours is [67], which shows that subtask decomposition enables learning the parity problem with polynomial sample complexity in recurrent neural networks (RNNs). Their learnability results rely on techniques from [57], which operate within the NTK linearization regime of RNNs. In contrast, the Transformers analyzed in our work exhibit feature learning and identify the sparse secret set within the attention module. On the positive side, a line of works that studies optimization dynamics of neural networks on parity [31, 5, 17, 15, 3, 4] show that $d^{\Omega(k)}$ samples is sufficient to learn parity, which is close to the statistical query lower bound. On the negative side, it has been shown that in SGD learning with additional noise, $d^{\Omega(k)}$ samples are indeed needed [52, 2]. However, it remains unclear whether standard mini-batch SGD will still need exponential sample complexity. In our paper, we take the alternative approach to consider the training as an online algorithm with bounded memory (parameters) and utilize results from online learning literature to provide a rigorous exponential lower bound [40, 30]. [21] shows that while parity is easy to represent using Transformers, the sensitivity structure of the function will require the representation to have large weight norms.
@@
Chain-of-Thoughts (CoT). Chain-of-Thoughts (CoT) is the technique to let the model generate a reasoning process before final answers. CoT prompting is an effective method in practice to improve language model’s reasoning capabilities [63, 73, 56, 76, 58, 14]. Training with CoT data has further improved the model’s capability in performing complex reasoning (see Qwen [49], Yue et al. [71], Yu et al. [70], Kim et al. [27] and reference therein). Different lines of work aimed at understanding the effect of CoT. From the representation theory perspective, works including [19, 42, 38, 44, 64] show that CoT can provably expand the representation power of different neural architectures including Transformers and RNNs. From the statistical approximation level, prompting with CoT has improved the statistical error [23, 47]. [35] studies how MLP models learn with CoT data. [33] studies how Transformers learn to perform CoT through prompting using gradient descent. We differ from this work as (1) we focus on how Transformer captures the sparse sequential dependency in Chain of Thought, which is not reflected in their data modeling; (2) we study zero-shot CoT data directly. Finally, Dutta et al. [16], Wang et al. [61] studies pretrained Transformers’ activation on CoT data. They highlight the attention head’s role in moving the essential information from the context to the reasoning process, which is consistent with our theoretical insight.
@@
Appendix A Omitted Proof
A.1 Notation and Assumptions.
@@ A.2 Representation Theory
We will set $d=\Theta(k\log(n/\delta))$ and $m=\Theta(k)$. We will use $p(x,y)$ as shorthand for $12(x-1)+y+$ 1.
@@
Now we only need to map from the summation $\sum_{j\in\mathcal{S}}\textbf{b}[j]$ to the parity of the summation. This can be done by choosing $a_{r}$ and $b_{r}$ iteratively. The proof is then complete. ∎
A.3 Dynamics without CoT
@@ A.4 Dynamics with CoT
    The secret Hamming weight satisfies $k\in\left[\frac{n}{\log^{3}(n/\delta)},\frac{n}{\log^{2}(n/\delta)}\right]$.
@@
    Define the batch size as $B=C_{2}n\log^{20}\left(\frac{n}{\delta}\right)$ for a sufficiently large constant $C_{2}$.
@@
The conditions in Lemma 5 are satisfied based on the definitions provided in Assumption LABEL:assum:order_concise. Below, we outline the verification for each condition:
@@
    $\frac{d}{n}=\frac{k\log^{1.1}\left(\frac{n}{\delta}\right)}{n}\in\left[\frac{1}{\log^{1.9}(n/\delta)},\frac{1}{\log^{0.9}(n/\delta)}\right]\to 0\text{ as }n\to\infty.$
@@
    $\frac{B}{nk}=\frac{C_{2}n\log^{20}\left(\frac{n}{\delta}\right)}{nk}=\frac{C_{2}\log^{20}\left(\frac{n}{\delta}\right)}{k}\leq\frac{C_{2}\log^{20}\left(\frac{n}{\delta}\right)}{\frac{n}{\log^{3}(n/\delta)}}=C_{2}\frac{\log^{23}(n/\delta)}{n}\to 0.$
@@
    $=\sqrt{\frac{2}{\log^{1.1}\left(\frac{n}{\delta}\right)}\cdot O(\log n)}=O\left(\frac{\sqrt{\log n}}{\log^{0.55}(n/\delta)}\right)\to 0.$
@@
      $\frac{\sqrt{2nk\log\left(\frac{300mnB}{\delta}\right)}}{\sqrt{Bd}}=\frac{\sqrt{2nk\cdot O(\log n)}}{\sqrt{C_{2}n\log^{20}\left(\frac{n}{\delta}\right)\cdot k\log^{1.1}\left(\frac{n}{\delta}\right)}}=O\left(\frac{\sqrt{nk\log n}}{\sqrt{nk}\cdot\log^{10.55}(n/\delta)}\right)=O\left(\frac{\sqrt{\log n}}{\log^{10.55}(n/\delta)}\right)\to 0.$
@@
    Given $k\geq\frac{n}{\log^{3}(n/\delta)}$,
    $\frac{n}{k\log^{11.65}(n/\delta)}\leq\frac{n}{\frac{n}{\log^{3}(n/\delta)}\log^{11.65}(n/\delta)}=\frac{1}{\log^{8.65}(n/\delta)}.$
@@
    $\log^{2}\left(\frac{300mnB}{\delta}\right)\frac{n\sqrt{nk}}{d\sqrt{Bd}}=O\left(\frac{\log^{2}(n/\delta)}{\log^{8.65}(n/\delta)}\right)=O\left(\log^{-6.65}(n/\delta)\right)\to 0\text{ as }n\to\infty.$
@@
    Substituting $C_{2}=\frac{1.28\times 10^{7}}{\epsilon^{2}}$:
    $\frac{\eta_{0}\eta_{1}}{256\,m\,n^{2}}=\frac{2k\times 1.28\times 10^{7}\log^{8}\left(\frac{n}{\delta}\right)}{256\times 10^{4}\times n}=10\times\frac{k\log^{8}\left(\frac{n}{\delta}\right)}{n}.$
    Using $k\geq\frac{n}{\log^{3}\left(\frac{n}{\delta}\right)}$ from the assumption:
    $\frac{\eta_{0}\eta_{1}}{256\,m\,n^{2}}\geq 10\times\frac{\frac{n}{\log^{3}\left(\frac{n}{\delta}\right)}\log^{8}\left(\frac{n}{\delta}\right)}{n}=10\log^{5}\left(\frac{n}{\delta}\right).$
    Since $\log^{5}\left(\frac{n}{\delta}\right)>\log n$ for sufficiently large $n$, we have:
    $10\log^{5}\left(\frac{n}{\delta}\right)>10\log n.$

### v2

@@ 1 Introduction
    Theoretically, we show that while the parity problem can be expressed by a 1-layer transformer, learning requires exponentially many samples without CoT when the number of parameters is limited (Theorem 1 and  2). Meanwhile, for CoT data with sparse sequential dependence, a 1-layer transformer can learn the parity function and faithfully represent the sparse dependence in its attention pattern with almost linear samples with respect to sequence length(Theorem 3).
@@ 3.1 Notations and Definitions
Parity Problem. Each token in our setting is either 00 or $1$. We represent a sequence of length $T$ as $\textbf{b}[1],\ldots,\textbf{b}[T]$. A parity function is a function of form $\mathrm{parity}_{S}(\textbf{b})=\oplus_{j\in\mathcal{S}}\textbf{b}[j]$, where $\mathcal{S}$ is the set of secret indices that is fixed during training and testing, and $\oplus$ is the XOR operator on binary variables. The cardinal of the secret set $k=|S|$ controls the hardness of the problem.
@@
Loss function. To simplify our analysis, we use hinge loss $\ell(\hat{y},y)=\max\{(-1)^{y}\hat{y}+1,0\}$ as the loss function. We define the next token prediction loss of a boolean sequence b as $L(w)=\ell(\mathcal{T}^{(L)}(\textbf{b})[n+1],\textbf{b}[n+2])$ and $L(w)=\sum_{i=n+1}^{n+k}\ell(\mathcal{T}^{(L)}(\textbf{b})[i],\textbf{b}[i+1])$ for with and without CoT setup respectively, where $w$ denotes all the trainable parameters.
@@ 3.2 Exponential Sample Complexity without CoT
Consider the Transformer model defined in Definition 6, for any $\delta<0.1$ and large enough $n$, when the number of secret indices $k$ is in $[n/\log^{5}(n/\delta),n/\log^{4}(n/\delta)]$, with probability at least $1-\delta$ over the randomness of embedding $e$, there exists a weight configuration of the Transformer with dimension $d=\Theta(k\log(n/\delta))$ and width $2m=O(k)$ with $\Theta(\log n)$ precision of the weights and activations, such that $d^{2}+dm=o(nk/\log n)$, and it achieves perfect accuracy on the parity problem $(n,k)$ without CoT, i.e.
$\forall\textbf{b}\sim p_{\rm{NoCoT}},\mathrm{sgn}\left(\mathcal{T}(\textbf{b})[n+1]\right)=(-1)^{\textbf{b}[n+2]+1}.$
In other words, the model possesses sufficient expressive power to represent any parity function, even in the absence of the chain of thought. Therefore, we are in the representational-sufficient regime where expressiveness is not the bottleneck. The proof of this statement is based on random matrix theory and concentration inequalities, which we defer to Section B.2.
@@
The proof is deferred to Section B.3. We use results from the classical online learning communities [43] to show that in the regime where the memory required to perform the training is less than $nk$, the model can’t effectively store information about enough samples inside the parameters during training, and hence can’t infer the secret indices effectively. Here the parameters $k$ denotes the size of the secret set $\mathcal{S}$ and the sample complexity grows exponentially when $k$ grows.
@@ 3.3 Polynomial Sample Complexity with CoT
We will use the following initialization of the 1-layer Transformer model. The attention is initialized to be uniform ($A=0$) and the contribution of attention output is initialized to be zero ($W_{r,d+1:2d}=0$). The parameter $h$ in the FFN and the word embedding $\mathbf{E}$ is fixed during training.
@@
Furthermore, after training, the attention pattern is interpretable and one-hot in the sense that, for any $\textbf{b}\sim p_{\rm{CoT}},i\in\{n+1,\cdots,n+k\},j\leq i$,
@@
This theorem indicates that the attention module can successfully extract the sparse sequential dependencies in the CoT data and faithfully represent it in the attention pattern, which is validated in our experiments (see Figure 3). It is also amongst the first optimization dynamics analysis of Transformers using finite-sample gradients rather than population gradients. We delay the full proof of the theorem to Section B.5.
@@
Phase 1. Configuring the FFN At initialization, as the FFN weight corresponds to the attention output ($W_{r,d+1:2d}$) is initialized to be zero. The set of neurons activated at each position $i\in[n+1,n+k]$ is solely determined by the word embedding $e_{(i,\textbf{b}[i])}$ at the position. Because $e_{i,b}$ are nearly orthogonal, it holds that $\langle e_{i,\mathbf{b}[i]},W_{r,1:d}\rangle\approx\nu_{r,i,b}$ for all $i$ and $r$ with high probability. As a result, for a fixed $i$ and $\mathbf{b}[i]$, the set of activated neurons at position $i$ is $\{r\mid\nu_{r,i,\textbf{b}[i]}>0\}$. Therefore, at initialization, the MLP can be viewed as an ensemble of multiple linear functions specialized to each position and boolean value. We can show that when the learning rate is small, the set of activated neurons at each position remains the same through the training process. Hence, we can conceptually view the FFN function as a set of different linear functions applied independently at each position and value. We will denote this set of linear weight as $\kappa_{t,i,b}$ (formally defined in Lemma 18).
@@ 4 Empirical Experiments
In this section, we validate and extend our theoretical findings through comprehensive experiments in the following three aspects: First, we empirically confirm that CoT reduces the sample complexity of transformers in learning the parity problem. Second, we conduct multi-pass training of transformers without CoT, as a complement to the established lower bound (Theorem 2) which only applies to constant-pass training. The results indicate that multi-pass training indeed improves the models’ ability to learn parity problems. Third, we conduct experiments on the GSM8K dataset [13] to show that CoT introduces sparse sequential dependence on real-world training data. Although necessary simplifications on transformer are made in Section 3 to develop theoretical results, we use the standard transformer architecture in the subsequent experiments following the GPT-2 architecture [53] with trainable position embeddings unless otherwise specified.
@@ 4.3 CoT Induces Sparsity on Real-World Data
Experiment protocols. In Figure 6, we examine two data types: With CoT, where inputs from GSM8K dataset are concatenated with ground truth answers that include multiple reasoning steps, and No CoT, where inputs are directly concatenated with the final answer. We evaluate two language models: the pre-trained model Qwen2-7B [71] and the specialized mathematics model Qwen2-7B-Math [52] which is fine-tuned from Qwen2-7B on a mathematics-specific corpus with CoT data. We plot the normalized attention entropy (Equation 1) across different heads. More details can be found in the Appendix Section C.1.
@@ 5 Additional Related Work
Parity Learning. The most relevant work to ours is [70], which shows that subtask decomposition enables learning the parity problem with polynomial sample complexity in recurrent neural networks (RNNs). Their learnability results rely on techniques from [60], which operate within the NTK linearization regime of RNNs. In contrast, the Transformers analyzed in our work exhibit feature learning and identify the sparse secret set within the attention module. On the positive side, a line of works that studies optimization dynamics of neural networks on parity [34, 6, 18, 16, 3, 5] show that $n^{\Omega(k)}$ samples is sufficient to learn parity, which is close to the statistical query lower bound [27]. On the negative side, it is well-established that learning parity using gradient descent (GD) requires $\Omega(n^{k})$ iterations due to the SQ lower bound [27]. For SGD learning, [55, 2] showed that an exponential number of samples is necessary when SGD is hindered by additional noise or when the number of weights updated in each step is constrained. However, it remains unclear whether standard mini-batch SGD will still need exponential sample complexity. In our paper, we take the alternative approach to consider the training as an online algorithm with bounded memory (parameters) and utilize results from online learning literature to provide a rigorous exponential lower bound [43, 33]. [22] shows that while parity is easy to represent using Transformers, the sensitivity structure of the function will require the representation to have large weight norms. A recent work [4] conjectured that a distribution is weakly learnable by a Transformer if and only if it has constant globality, which is supported by empirical experiments. The work also highlights the parity problem as a notable special case.
@@
Chain-of-Thoughts (CoT). CoT is the technique to let the model generate a reasoning process before final answers. CoT prompting has proven effective in enhancing language models’ reasoning capabilities [66, 76, 59, 79, 61, 15]. Training with CoT data has further improved the model’s capability in performing complex reasoning (see Qwen [52], Yue et al. [74], Yu et al. [73], Kim et al. [30] and reference therein). Different lines of work have explored the effect of CoT. From the representation theory perspective, works including [20, 45, 41, 47, 67] show that CoT can provably expand the representation power of neural architectures including Transformers and RNNs. From the statistical approximation level, prompting with CoT reduces the statistical error [24, 50]. [38] studies how MLP learns with CoT data. [36] studies how Transformers learn to perform CoT through prompting using gradient descent. We differ from this work as (1) we focus on how Transformer captures the sparse sequential dependency in CoT, which is not reflected in their data modeling; (2) we study zero-shot CoT data directly. Dutta et al. [17], Wang et al. [64] studies pre-trained Transformers’ activation on CoT data. They highlight the attention head’s role in moving the essential information from the context to the reasoning process, which is consistent with our theoretical insight. Concurrently with our work, Kim and Suzuki [28] theoretically demonstrated that CoT reduces the number of gradient steps required to solve the parity problem. While their analysis focuses on GD with a noisy oracle, our work investigates mini-batch SGD and provides theoretical guarantees on the exponential reduction in sample complexity.
@@
Acknowledgements
Jingzhao Zhang is supported by National Key R&D Program of China 2024YFA1015800 and Shanghai Qi Zhi Institute Innovation Program.
Appendix A Additional Related Work
Appendix B Omitted Proof
B.1 Notation and Assumptions.
@@ B.2 Representation Theory
We will set $d=\Theta(k\log(n/\delta))$ and $m=k+1$. We will use $p(x,y)$ as shorthand for $12(x-1)+y+$ 1.
@@
If we define
$z_{j}=e_{j,\textbf{b}[j]}^{T}Ae_{n+1,0}.$
The softmax value is given by:
$\mathrm{softmax}(z)_{j,n+1}=\frac{e^{z_{j}}}{\sum_{j^{\prime}}e^{z_{j^{\prime}}}}.$
  1. 1.
    For $j\in\mathcal{S}$:
    $e^{z_{j}}=e^{40\log n}=n^{40}.$
  2. 2.
    For $j\not\in\mathcal{S}$:
    $e^{z_{j}}\leq e^{20\log n}=n^{20}.$
With $|\mathcal{S}|=k$, the denominator of the softmax becomes:
$\sum_{j^{\prime}}e^{z_{j^{\prime}}}=\sum_{j^{\prime}\in\mathcal{S}}e^{40\log n}+\sum_{j^{\prime}\notin\mathcal{S}}e^{z_{j^{\prime}}}\in[kn^{40},kn^{40}+n^{21}].$
For $j\in\mathcal{S}$:
$\mathrm{softmax}(z)_{j,n+1}=\frac{n^{40}}{kn^{40}+O(n^{20})}=\frac{1}{k}\cdot\frac{1}{1+O(n^{-19})}=\frac{1}{k}+O(n^{-19}).$
For $j\notin\mathcal{S}$:
$\mathrm{softmax}(z)_{j,n+1}=\frac{e^{z_{j}}}{kn^{40}+O(n^{20})}\leq\frac{n^{20}}{kn^{40}}=O(n^{-20}).$
@@
Now we only need to map from the summation $\sum_{j\in\mathcal{S}}\textbf{b}[j]$ to the parity of the summation.We set
$b_{r}$ $=2k,$
$a_{r}$ =$=\left\{\begin{array}[]{ll}-4\lceil\frac{r}{2}\rceil+4&\text{if }1\leq r\leq m,\\ 1&\text{if }r=m+1,\\ -4\lfloor\frac{r-m}{2}\rfloor+2&\text{if }m+2\leq r\leq 2m.\end{array}\right.$
$h_{r}$ =$=\left\{\begin{array}[]{ll}1&\text{if }1\leq r\leq m,\\ -1&\text{if }m+1\leq r\leq 2m,\\ \end{array}\right.$
Then
$\mathcal{T}(\mathbb{\mathbf{E}}(\textbf{b}))[n+1]$ $=\sum_{r=1}^{k}h_{r}\mathrm{ReLU}\left(a_{r}-b_{r}\langle U,\mathcal{A}(\mathbf{E}(\textbf{b}))[n+1]\rangle\right)$
$=\sum_{r=1}^{m}\mathrm{ReLU}(a_{r}+2\sum_{j\in S}\textbf{b}[j])-\sum_{r=m}^{2m}\mathrm{ReLU}(a_{r}+2\sum_{j\in S}\textbf{b}[j])+O(\frac{1}{n^{4}})$
$=(-1)^{\sum_{j\in S}\textbf{b}[j]+1}+O(\frac{1}{n^{4}})$
$=(-1)^{\textbf{b}[n+2]+1}+O(\frac{1}{n^{4}}).$
Thus $\mathrm{sgn}(\mathcal{T}(\mathbb{\mathbf{E}}(\textbf{b}))[n+1])=(-1)^{\textbf{b}[n+2]}$ for large enough $n$.
Note that the range of parameters is polynomial with $n$, thus could be represented with $\Theta(\log n)$ precision. With $\Theta(\log n)$ precision, the error of each activation can be bounded by a small inverse polynomial of $n$ and hence we can show the same result.
B.3 Dynamics without CoT
@@ B.4 Dynamics with CoT
    The secret Hamming weight satisfies $k\in\left[\frac{n}{\log^{5}(n/\delta)},\frac{n}{\log^{4}(n/\delta)}\right]$.
@@
    Define the batch size as $B=C_{2}n\log^{20}\left(\frac{n}{\delta}\right)$ for a sufficiently large constant $C_{2}=\frac{1.28\times 10^{7}}{\epsilon^{2}}$.
@@
The conditions in Lemma 5 are satisfied based on the definitions provided in Assumption 2. Below, we outline the verification for each condition:
@@
    $\frac{d}{n}=\frac{k\log^{1.1}\left(\frac{n}{\delta}\right)}{n}\in\left[\frac{1}{\log^{3.9}(n/\delta)},\frac{1}{\log^{2.9}(n/\delta)}\right]\to 0\text{ as }n\to\infty.$
@@
    $\frac{B}{nk}=\frac{C_{2}n\log^{20}\left(\frac{n}{\delta}\right)}{nk}=\frac{C_{2}\log^{20}\left(\frac{n}{\delta}\right)}{k}\leq\frac{C_{2}\log^{20}\left(\frac{n}{\delta}\right)}{\frac{n}{\log^{5}(n/\delta)}}=C_{2}\frac{\log^{25}(n/\delta)}{n}\to 0.$
@@
    $\sqrt{\frac{2k}{d}\log\left(\frac{Bmn}{\delta}\right)}=\sqrt{\frac{2}{\log^{1.1}\left(\frac{n}{\delta}\right)}\cdot O(\log n)}=O\left(\frac{\sqrt{\log n}}{\log^{0.55}(n/\delta)}\right)\to 0.$
@@
      $\frac{\sqrt{2nk\log\left(\frac{300mnB}{\delta}\right)}}{\sqrt{Bd}}=\frac{\sqrt{2nk\cdot O(\log n)}}{\sqrt{C_{2}n\log^{20}\left(\frac{n}{\delta}\right)\cdot k\log^{1.1}\left(\frac{n}{\delta}\right)}}=O\left(\frac{\sqrt{nk\log n}}{\sqrt{nk}\cdot\log^{10.55}(n/\delta)}\right)$
      $=O\left(\frac{\sqrt{\log n}}{\log^{10.55}(n/\delta)}\right)\to 0.$
@@
    Given $k\geq\frac{n}{\log^{5}(n/\delta)}$,
    $\frac{n}{k\log^{11.65}(n/\delta)}\leq\frac{n}{\frac{n}{\log^{5}(n/\delta)}\log^{11.65}(n/\delta)}=\frac{1}{\log^{5.65}(n/\delta)}.$
@@
    $\log^{2}\left(\frac{300mnB}{\delta}\right)\frac{n\sqrt{nk}}{d\sqrt{Bd}}=O\left(\frac{\log^{2}(n/\delta)}{\log^{5.65}(n/\delta)}\right)=O\left(\log^{-4.65}(n/\delta)\right)\to 0\text{ as }n\to\infty.$
@@
    Setting $C_{2}=\frac{1.28\times 10^{7}}{\epsilon^{2}}$:
    $\frac{\eta_{0}\eta_{1}}{256\,m\,n^{2}}\geq\frac{2k\times 1.28\times 10^{7}\log^{8}\left(\frac{n}{\delta}\right)}{256\times 10^{4}\times n}=10\times\frac{k\log^{8}\left(\frac{n}{\delta}\right)}{n}.$
    Using $k\geq\frac{n}{\log^{5}\left(\frac{n}{\delta}\right)}$ from the assumption:
    $\frac{\eta_{0}\eta_{1}}{256\,m\,n^{2}}\geq 10\times\frac{\frac{n}{\log^{5}\left(\frac{n}{\delta}\right)}\log^{8}\left(\frac{n}{\delta}\right)}{n}=10\log^{3}\left(\frac{n}{\delta}\right).$
    Since $\log^{3}\left(\frac{n}{\delta}\right)>4\log n$ for sufficiently large $n$, we have:
    $10\log^{3}\left(\frac{n}{\delta}\right)>40\log n.$